### v1

@@ Abstract
When learning stable linear dynamical systems from data, three important properties are desirable: i) predictive accuracy, ii) provable stability, and iii) computational efficiency. Unconstrained minimization of reconstruction errors leads to high accuracy and efficiency but cannot guarantee stability. Existing methods to remedy this focus on enforcing stability while also ensuring accuracy, but do so only at the cost of increased computation. In this work, we investigate if a straightforward approach can simultaneously offer all three desiderata of learning stable linear systems. Specifically, we consider a post-hoc approach that manipulates the spectrum of the learned system matrix after it is learned in an unconstrained fashion. We call this approach spectrum clipping (SC) as it involves eigen decomposition and subsequent reconstruction of the system matrix after clipping all of its eigenvalues that are larger than one to one (without altering the eigenvectors). Through detailed experiments involving two different applications and publicly available benchmark datasets, we demonstrate that this simple technique can simultaneously learn highly accurate linear systems that are provably stable. Notably, we demonstrate that SC can achieve similar or better performance than strong baselines while being orders-of-magnitude faster. We also show that SC can be readily combined with Koopman operators to learn stable nonlinear dynamics, such as those underlying complex dexterous manipulation skills involving multi-fingered robotic hands. Our codes and dataset can be found at https://github.com/GT-STAR-Lab/spec_clip.
@@ 1 Introduction
A number of methods have been developed to enforce stability of learned linear systems and reduce sensitivity to training data (Lacy and Bernstein, 2002, 2003; Boots et al., 2007; Huang et al., 2016; Gillis et al., 2020; Mamakoukas et al., 2020, 2023). These methods take one of three main approaches: a) minimizing the regression objective under the stability constraints (Lacy and Bernstein, 2002, 2003), b) iterating between solving an unconstrained problem and optimizing its solution under the stability constraints (Boots et al., 2007; Huang et al., 2016), and c). iterating between characterizing the stable matrix and optimizing the characterizations with the regression objective (Gillis et al., 2020; Mamakoukas et al., 2020, 2023). Regardless of the underlying approach, these methods inevitably increase the computation burden, potentially counteracting one of the primary advantages of leveraging linear techniques. Furthermore, they exhibit poor scalability when applied to high-dimensional systems, which are necessary in both modeling inherently high-dimensional state spaces (e.g., images) and effectively approximating complex nonlinear systems using Koopman operators (e.g., robotic systems).
In this work, we suggest and investigate a simple yet surprisingly effective method, Spectrum Clipping (SC), to learn the stable linear systems from data requiring negligible additional computations. This approach involves a straightforward implementation: 1) compute the system evolution matrix by solving the least-squares regression problem, 2) perform eigen decomposition of the system matrix and extract all eigenvalues, 3) clip eigenvalues with magnitudes larger than one to one, and 4) reconstruct the system evolution matrix using the clipped eigenvalues and the unchanged eigenvectors. SC enforces stability by construction but neither requires complex constrained optimization nor updating the matrix iteratively. As a result, SC is both efficient and scales well to high-dimensional systems. We are inspired to investigate this simple approach since strategies related to clipping a matrix’s spectrum have proven effective in other domains. For instance, directly controlling the singular values of linear layers seems to enhance training stability and reduce the model’s generalization error (Boroojeny et al., 2024; Sedghi et al., 2018; Senderovich et al., 2022).
@@
We also investigated if SC can be useful in learning complex nonlinear systems when combined with Koopman operators. Specifically, we evaluated SC’s ability to learn stable linear systems in the lifted space that encode complex dexterous manipulation skills for a multi-fingered robotic hand (Han et al., 2023). Our analyses reveals that unconstrained learning produces unstable linear systems if the training data includes unsuccessful demonstrations or when the task horizon is truncated. We are motivated by the fact that such characteristics are commonly observed when robot learning data is collected from human (Qin et al., 2022; Shaw et al., 2023; Bahl et al., 2022). Our results show that SC can be readily combined with Koopman operators to effectively learn stable nonlinear systems in mere seconds, while remaining robust to the undesirable characteristics in the training data.

### v3

@@ Abstract
When learning stable linear dynamical systems from data, three important properties are desirable: i) predictive accuracy, ii) provable stability, and iii) computational efficiency. Unconstrained minimization of reconstruction errors leads to high accuracy and efficiency but cannot guarantee stability. Existing methods to remedy this focus on enforcing stability while also ensuring accuracy, but do so only at the cost of increased computation. In this work, we investigate if a straightforward approach can simultaneously offer all three desiderata of learning stable linear systems. Specifically, we consider a post-hoc approach that manipulates the spectrum of the learned system matrix after it is learned in an unconstrained fashion. We call this approach spectrum clipping (SC) as it involves eigen decomposition and subsequent reconstruction of the system matrix after clipping all of its eigenvalues that are larger than one to one (without altering the eigenvectors). Through detailed experiments involving two different applications and publicly available benchmark datasets, we demonstrate that this simple technique can simultaneously learn highly accurate linear systems that are provably stable. Notably, we demonstrate that SC can achieve similar or better performance than strong baselines while being orders-of-magnitude faster. We also show that SC can be readily combined with Koopman operators to learn stable nonlinear dynamics, such as those underlying complex dexterous manipulation skills involving multi-fingered robotic hands. Further, we find that SC can learn stable robot policies even when the training data includes unsuccessful or truncated demonstrations. Our codes and dataset can be found at https://github.com/GT-STAR-Lab/spec_clip.
@@ 1 Introduction
A number of methods have been developed to enforce stability of learned linear systems and reduce sensitivity to training data (Lacy and Bernstein, 2002, 2003; Boots et al., 2007; Huang et al., 2016; Gillis et al., 2020; Mamakoukas et al., 2020, 2023). These methods take one of three main approaches: a) minimizing the regression objective under the stability constraints (Lacy and Bernstein, 2002, 2003), b) iterating between solving an unconstrained problem and optimizing its solution under the stability constraints (Boots et al., 2007; Huang et al., 2016), and c). iterating between characterizing the stable matrix and optimizing the characterizations with the regression objective (Gillis et al., 2020; Mamakoukas et al., 2020, 2023). Regardless of the underlying approach, these methods inevitably increase the computation burden, potentially counteracting one of the primary advantages of leveraging linear techniques. Furthermore, they exhibit poor scalability when applied to high-dimensional systems, which are necessary in both modeling inherently high-dimensional state spaces (e.g., images) and effectively approximating complex nonlinear systems using Koopman operators Koopman (1931) (e.g., robotic systems).
In this work, we suggest and investigate a simple yet surprisingly effective method, Spectrum Clipping (SC), to learn the stable linear systems from data requiring negligible additional computations. This approach involves a straightforward implementation: 1) compute the system evolution matrix by solving the least-squares regression problem, 2) perform eigen decomposition of the system matrix and extract all eigenvalues, 3) clip eigenvalues with magnitudes larger than one to one, and 4) reconstruct the system evolution matrix using the clipped eigenvalues and the unchanged eigenvectors. SC enforces stability by construction but neither requires complex constrained optimization nor updating the matrix iteratively. As a result, SC is both efficient and scales well to high-dimensional systems. We are inspired to investigate this simple approach since strategies related to clipping a matrix’s spectrum have proven effective in other domains. For instance, directly controlling the singular values of linear layers seems to enhance training stability and reduce the model’s generalization error (Boroojeny et al., 2024; Sedghi et al., 2018; Senderovich et al., 2022). We are also inspired by seminal works that revealed the spectral properties of the Koopman operator (Rowley et al., 2009; Mezić, 2013; Mezic, 2020). Indeed, if one clips the spectrum of a learned Koopman matrix, this simple approach can help enforce stability when learning nonlinear systems.
@@
We also investigated if SC can be combined with Koopman operators to learn complex nonlinear systems. Specifically, we evaluated SC’s ability to learn stable linear systems in the lifted space that encode complex dexterous manipulation skills for a multi-fingered robotic hand (Han et al., 2023). Our analyses reveals that unconstrained learning produces unstable linear systems if the training data includes unsuccessful demonstrations or when the task horizon is truncated. We are motivated by the fact that such characteristics are commonly observed when robot learning data is collected from humans (Qin et al., 2022; Shaw et al., 2023; Bahl et al., 2022). Our results show that SC can be readily combined with Koopman operators to learn stable nonlinear systems in mere seconds, while remaining robust to the undesirable characteristics in the training data. We also find that unstable and erratic robot behavior can arise due to a subset of the underlying dynamic modes (see Schmid (2022)). We find that clipping the spectrum of the learned Koopman matrix effectively stabilizes its rollouts and enables the robot to safely complete the task. These findings are in line with the well-established connections between Koopman modes and the state-space geometry of the underlying dynamical systems (Rowley et al., 2009; Mezić, 2013; Mezic, 2020).
@@ 4.6 SC can learn stable LDSs from shorter and corrupted datasets
We also reveal the connections between SC and the Koopman Operator’s spectral properties. Following the algorithm described in Rowley et al. (2009), we can rollout the systems with only subsets of the Koopman modes and visualize the hand motions to investigate the effectiveness of SC upon each subset.
Specifically, for each eigenvalue $\lambda_{i}\ (i=1,...,n)$ of system matrix $K$, we can find corresponding eigenvector $\mathbf{v_{i}}$ (i.e., the mode) and eigenvector $\mathbf{w_{i}}$ of adjoint matrix $K^{*}$, such that $K\mathbf{v_{i}}=\lambda_{i}\mathbf{v_{i}}\text{, and }K^{*}\mathbf{w_{i}}=\overline{\lambda_{i}}\mathbf{w_{i}}$. Then following (Rowley et al., 2009), we construct the corresponding eigenfunction $\varphi_{i}:\mathbb{R}^{n}\rightarrow C$, defined as $\varphi_{i}(\mathbf{z})=\langle\mathbf{z},\mathbf{w}_{i}\rangle$. By doing so, the rollout of eigenfunctions can be represented by $\varphi_{i}(K\mathbf{z})=\lambda_{i}\varphi_{i}(\mathbf{z})$, and the states can be reconstructed by $\mathbf{z_{k}}=\sum_{i=1}^{n}\lambda_{i}^{k}\varphi_{i}(\mathbf{z_{0}})\mathbf{v_{i}}$. Based on this formula, specific modes can be selected for rollouts.
In our experiments, we first computed the Koopman matrix $K_{f}$ from $Reloc\_with\_failures$ dataset. We then sorted the eigenvalues by their norms from largest to smallest and divided them into two subsets for rollouts: (i) Unstable Set: the first 52 Koopman modes with unstable eigenvalues, which have norms greater than 1, and (ii) Stable Set: the remaining 707 modes. We then visualized the robot hand motions of each subset via the eigenfunction rollouts in Fig. 8, respectively. It can be seen that the rollout of the Unstable Set before clipping clearly generate unstable hand motions (see upper line of the green box), which results in the low success rate and safety rate when combined with Stable Set for the final task performance (as shown in Table. 3). Further, we clipped the matrix $K_{f}$ and visualized the rollouts of the clipped Unstable Set in Fig. 8. It is obvious that after clipping, the robot hand motions become stable (see bottom line of the green box), enabling the composite system to safely complete the task (as also shown in Table 3).
Figure 8: Rollout with two subsets of the Koopman modes. The green box represents rollouts using Unstable Set (norm of eigenvalue larger than 1). The blue box represents rollout with the rest Stable Set. In the green box, the upper line shows the rollout before clipping and the lower line shows it after clipping. It is obvious that after clipping, the rollout becomes stable, thereby allowing the composite rollout of both sets to safely complete the task (as shown in Table 3).